OpenStudy (anonymous):

Describe the graph of this inequality: 3x + 2y < 4

OpenStudy (anonymous):

So you first graph 3x+2y=4. this can be rewritten as y=-3/2 x+2 you see that the y-intercept is 2 and that the slope is -3/2 so you have  but you have a broken line b/c of the "<" so now you pick any point in the plane...let's try (0,0) when x=0 and y=0 we have 3(0)+2(0)=0 which is <4, so (0,0) makes our inequality statement 3x+2y<4 true so the side that has (0,0) is shaded in.